# ContrastAlign: Toward Robust BEV Feature Alignment via Contrastive Learning for Multi-Modal 3D Object Detection

### v2

@@ 4.1.2 Implementation Details
TransFusion [1] LC 66.4 70.6 86.3 61.8 26.9 75.1 42.0 73.1 74.9 62.5 85.2 75.9
Noisy BEVFusion [34] LC 60.8 65.7 83.1 50.3 26.5 66.4 38.0 65.0 64.9 52.8 86.1 75.1
ContrastAlign LC 68.1 70.9 88.6 63.9 29.3 74.5 41.8 71.6 77.9 64.9 88.6 80.0
+7.3 +5.2 +5.5 +13.6 +2.8 +8.1 +3.8 +6.6 +13.0 +12.1 +2.5 +4.9
Table 2: Comparison with the SOTA methods on the nuScenes test set. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. ‘L’ means only LiDAR data are used, ‘LC’ denotes the use of both LiDAR and camera data. means using TTA (test-time augmentation). The best performances are marked in bold.
@@
ContrastAlign LC 71.8 73.8 89.0 60.9 41.1 70.1 64.6 82.2 75.9 53.8 90.9 89.3
+1.6 +0.9 +0.4 +0.8 +1.8 +0.3 +0.8 +2.2 +1.8 +2.8 +1.7 +2.8
4.2 Comparison with the SOTA Methods
4.2.1 Results on Clean and Noisy Misalignment Settings
As shown in Table 1, we compare our ContrastAlign with SOTA methods on the nuScenes validation set under clean and noisy misalignment settings. In the clean nuScenes validation set, our method achieves SOTA performance 70.3% mAP and 72.5% NDS, outperforming the baseline (BEVFusion [34]) by 1.8% mAP and 1.1% NDS. Compared to SOTA methods include TransFusion-L [1], FUTR3D [6], TransFusion [1], and ObjectFusion [4], our method achieves the best performance, especially in challenging categories like construction vehicles, trailers, barriers, bikes, and traffic cones, which mainly consist of small and difficult-to-detect objects.
Our main focus is not on the clean benchmark but on addressing the noisy misalignment issue. In comparison to TransFusion [1] and BEVFusion [34] on the nuScenes validation set under noisy misalignment setting, our ContrastAlign achieves SOTA performance. Compared to the baseline (BEVFusion [34]), our ContrastAlign improves performance by 7.3% mAP and 5.2% NDS, with particularly significant improvements of 13.6%, 13.0%, and 12.1% in the categories of truck, motorcycle, and bike, respectively. It is worth noting that TransFusion [1], as a feature-level method, addresses feature misalignment through attention. However, the global query scheme of TransFusion cannot be directly applied to BEVFusion [34, 29] due to differences in their architectures or objectives. In summary, our ContrastAlign has shown significant improvement not only in the clean setting but also in the noise setting, proving that the proposed method enables to deal with the misalignment issue.
4.2.2 Results on NuScenes Test Set
As shown in Table 2, ContrastAlign achieves 71.8% mAP and 73.8% NDS on the nuScenes test benchmark, outperforming BEVFusion [34] by 1.6% mAP and 0.9% NDS. Compared to both LiDAR-based methods such as CenterPoint [66], VoxelNeXt [9], TransFusion-L [1], and multi-modal methods such as TransFusion [1], DeepInteraction [64], and ObjectFusion [4], our method achieves SOTA performance. Specifically, compared to ObjectFusion [4], as a solution to BEV feature misalignment, our method outperforms ObjectFusion by 0.8% mAP and 0.5% NDS. What’s more, ContrastAlign has significant improvements of 1.8%, 2.2%, 1.8%, 2.8%, 1.7%, and 2.8% in construction vehicles, barriers, motorcycles, bikes, pedestrians, and traffic cones, respectively. Although the nuScenes is an open-source clean dataset, it inevitably has minimal feature misalignment issues. Our notable improvements on the nuScenes test set indicate that our method effectively mitigates feature misalignment.
4.3 Robustness Analysis
To demonstrate the robustness of our method, we present the results in Table 3 under different times, ego distances, and object sizes. Since different times pose challenges to the detection performance of the model, we follow the procedure of BEVFusion [34] and split the validation set into Day/Night based on the keywords ‘day’ and ‘night’ in the scene descriptions. Compared to BEVFusion [34], our method significantly improves Day and Night scenarios through contrastive learning for feature alignment. Specifically, for night scenarios, our method outperforms BEVFusion by 1.3%, where depth estimation presents a greater challenge for BEVFusion due to poor lighting conditions.
Far-small objects are more susceptible to feature misalignment. Following BEVFusion [34], we categorize annotation and prediction ego distances into three groups, Near (0-20m), Middle (20-30m) and Far (>30m), summarize the size distributions for each category, defining three equal-proportion size levels, Small, Moderate and Large. For the different ego distance, we have seen improvements in Near, Middle, and Far, especially in Far, which has increased by 3.9%. In the different object sizes, we have seen significant improvements in Small, Moderate, and Large, especially in Small objects, which has increased by 3.1%. The improvement of the ‘Far’ and ‘Small’ metrics emphasizes the improvement of our method, which makes our method be visually meaningful since better feature alignment allows for better detection of smaller and farther objects.
@@ 4.3 Robustness Analysis
Table 3: Robustness to weather conditions, different ego distances, different sizes on nuScenes [3] clean validation set. The evaluation metric is mAP (%).
@@
Table 4: Roles of different modules in ContrastAlign for feature alignment on nuScenes validation set under noisy misalignment setting. ‘C.V.’, ‘Motor.’, ‘Ped.’, ‘T.C.’, ‘C.L.’, ‘I.F.’, and ‘LT (ms)’ are short for construction vehicle, motorcycle, pedestrian, traffic cone, C-Instance and L-Instance modules, the InstanceFusion module, and latency respectively. All latency measurements are conducted on the same workstation with an A100 GPU.
[34] C.L. I.F. mAP NDS LT (ms) Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
@@
Table 5: Effect of $\tau$ in InstanceFusion module.
$\tau$ mAP NDS LT (ms)
0.05 67.3 69.2 156.3
0.1 68.1 70.9 154.4
0.2 65.0 67.3 151.2
Table 6: Effect of $K$ in InstanceFusion module.
$K$ mAP NDS LT (ms)
5 67.8 69.7 151.3
8 68.1 70.9 154.4
16 67.2 68.1 159.8
4.4 Ablation Studies
@@
To analyze the impact of different modules in ContrastAlign on feature misalignment, we conduct experiments on nuScenes validation set under the noise misalignment setting, as shown in Table 4. It can be seen that after adding InstanceFusion module, there is a significant improvement, with mAP increasing from 61.2% to 68.1% (an increase of 6.9%), which is a very significant performance improvement. Meanwhile, the impact on latency is within a controllable range. Evidence demonstrates the effectiveness of incorporating contrastive learning into multi-modal 3D object detection to address feature misalignment issues in BEV features.
@@
Furthermore, to assess the influence of various hyperparameters within the InstanceFusion module, an examination of the variables $\tau$ and $K$ was conducted on nuScenes validation set under noisy misalignment setting. Where $\tau$ comes from Eq.(1), which is the threshold of IoU affecting the number of positive sample pairs, and $K$ comes from Eq.(3), affecting the number of negative samples. Due to the different positive and negative sample pairs at different frames, we cannot provide a fixed number. As shown in Table 5, when $\tau$ changes, $K$ is fixed at 8 to obtain better performance indicators. As shown in Table 6, when $K$ changes, $\tau$ is fixed at 0.1. We choose several representative results for analysis. The larger the $\tau$, the smaller the number of positive samples. It can be observed that when $\tau$ is 0.1, the performance is the best, reaching 68.1% mAP and 70.9% NDS. The larger $K$ , the larger the number of negative samples, and the best performance is achieved when $K$ is 8.
5 Conclusion

### v3

@@ V Experiments
GraphBEV [38] LC 70.1 72.9 89.9 64.7 31.1 76.0 43.8 76.0 80.1 67.5 89.2 82.2
ContrastAlign LC 71.5 73.7 90.1 67.1 33.1 77.2 46.4 77.5 81.2 68.4 90.6 83.5
+1.4 +0.8 +0.2 +2.4 +2.0 +1.2 +2.6 +1.5 +1.1 +0.9 +1.4 +1.3
TABLE II: Comparison with SOTAs on the nuScenes val dataset with nosiy setting [52] including Spatial Misalignment & Temporal Misalignment. Blue denotes clean mAP, Red denotes noisy mAP, and RCE denotes Relative Corruption Error from Ref.[52].
Method mAPclean mAPnoisy RCE ↓ (%)
Spatial Temporal Spatial Temporal
Focals Conv-F [55] 67.7 18.7 21.3 72.4 68.5
AutoAlignV2 [31] 68.5 62.9 62.8 8.2 8.3
TransFusion [28] 66.4 59.2 43.7 10.8 34.2
FUTR3D [72] 64.2 58.8 51.4 8.4 19.9
DeepInteraction [30] 69.9 62.1 59.4 11.1 15.0
CMT [73] 70.3 65.2 58.7 7.2 16.5
BEVFusion [1] 68.5 60.8 49.0 11.2 28.5
ContrastAlign 70.3 68.1 63.1 3.1 10.2
GraphBEV [38] 70.1 69.1 55.1 1.4 21.4
ContrastAlign 71.5 70.5 66.2 1.4 7.4
TABLE III: Comparison with the SOTA methods on the nuScenes test set. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. ‘L’ means only LiDAR data are used, ‘LC’ denotes the use of both LiDAR and camera data. The best performances are marked in bold.
@@
ContrastAlign LC 72.2 73.8 89.0 60.9 41.1 73.1 64.6 82.2 75.9 54.5 90.9 89.3
+2.0 +0.9 +0.4 +0.8 +1.8 +4.3 +0.8 +2.2 +1.8 +3.5 +1.7 +2.8
GraphBEV [38] LC 71.7 73.6 89.2 60.0 40.8 72.1 64.5 80.1 76.8 53.3 90.9 88.9
ContrastAlign LC 72.9 74.9 89.5 61.5 41.8 74.5 65.7 82.6 77.4 54.8 91.6 90.0
+1.2 +1.3 +0.3 +1.5 +1.0 +2.4 +1.2 +2.5 +0.6 +1.5 +0.7 +1.1
TABLE IV: Comparison with prior methods on Argoverse2 validation set. Metrics: mAP (%)↑ for the overall results, AP (%)↑ for each category. * denotes result from Ref. [68]. denotes result re-implentment.
Method mAP Veh. Bus Ped. Stop. Box. Boll. C-B. M.-list MPC. M.-cycle Bicycle A-B. School. Truck. C-C. V-T. Sign Large. Str. Bic.-list
CenterPoint* [67] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1
FSD* [79] 28.2 68.1 40.9 59.0 29.0 38.5 41.8 42.6 39.7 26.2 49.0 38.6 20.4 30.5 14.8 41.2 26.9 11.9 5.9 13.8 33.4
VoxelNeXt* [68] 30.0 71.7 39.2 63.1 39.2 40.0 52.5 63.7 42.2 34.9 42.7 40.1 20.1 25.2 16.9 45.7 22.3 15.8 5.9 9.8 33.5
BEVFusion [1] 43.1 83.1 48.1 64.2 48.2 54.5 58.1 65.3 45.1 39.8 49.3 48.1 33.1 38.1 28.5 54.1 37.4 34.6 31.2 34.4 38.1
ContrastAlign 45.7 85.0 51.1 67.0 51.7 57.0 61.2 68.0 48.0 42.5 53.5 52.0 35.5 42.3 31.2 57.0 41.0 36.5 35.3 36.8 41.2
GraphBEV [38] 45.1 84.7 51.6 66.5 51.9 57.4 60.8 67.2 48.3 43.0 53.1 52.5 35.2 42.0 30.8 56.8 41.3 36.0 35.5 36.6 41.0
ContrastAlign 47.1 86.0 52.4 68.5 53.2 59.7 62.0 69.5 49.0 43.6 54.0 53.5 37.5 43.5 32.0 58.0 41.5 38.0 37.0 37.5 42.0
@@
TABLE V: Robustness to weather conditions, different ego distances, different sizes on nuScenes [51] clean validation set. The evaluation metric is mAP (%).
@@
GraphBEV [38] 69.7 45.1 78.6 65.3 42.1 55.4 58.3 63.1
ContrastAlign 70.3+0.6 47.0 +1.9 79.9 +1.3 66.1 +0.8 44.6 +1.5 58.3 +2.9 60.2 +1.9 64.6 +1.5
TABLE VI: Roles of different modules in ContrastAlign for feature alignment on nuScenes validation set under noisy misalignment setting. ‘C.V.’, ‘Motor.’, ‘Ped.’, ‘T.C.’, ‘C.L.’, ‘I.F.’, and ‘LT (ms)’ are short for construction vehicle, motorcycle, pedestrian, traffic cone, C-Instance and L-Instance modules, the InstanceFusion module, and latency respectively. All latency measurements are conducted on the same workstation with an A100 GPU.
BEVFusion C.L. I.F. mAP NDS LT (ms) Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
@@
TABLE VII: The impact of hyperparameter $\tau$ in InstanceFusion module.
$\tau$ mAP NDS LT (ms) Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
0.05 67.3 69.2 156.3 88.1 63.5 29.8 74.0 41.3 71.1 77.4 64.4 88.1 79.5
0.1 68.1 70.9 154.4 88.6 63.9 29.3 74.5 41.8 71.6 77.9 64.9 88.6 80.0
0.2 65.0 67.3 151.2 84.7 60.5 24.6 71.8 39.8 69.4 75.2 62.5 85.0 76.6
TABLE VIII: The impact of hyperparameter $K$ in InstanceFusion module.
$K$ mAP NDS LT (ms) Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
5 67.8 69.7 151.3 87.5 62.5 28.4 73.5 40.6 70.3 75.9 63.2 86.9 78.9
8 68.1 70.9 154.4 88.6 63.9 29.3 74.5 41.8 71.6 77.9 64.9 88.6 80.0
16 67.2 68.1 159.8 87.1 61.9 28.0 73.2 39.9 69.8 75.5 62.6 85.8 78.2
V-A Experimental Setup
V-A1 NuScenes Dataset and Evaluation Metrics
We evaluate the 3D object detection performance of the proposed ContrastAlign by comparing it with other state-of-the-art approaches on the nuScenes benchmark [51], which is collected with a 32-beam LiDAR and 6 cameras. The six images cover a 360-degree surrounding, and the dataset provides calibration matrices that enable precise projection from 3D points to 2D pixels. It requires detecting 10 object categories that are commonly observed in driving scenarios. We use mAP and NDS across all categories as the primary metrics for evaluation following [28, 1, 2, 38]. Note that the NDS metric is a weighted average of mAP and other breakdown metrics (e.g., translation, scale, orientation, velocity, and attributes errors).
V-A2 Argoverse 2 (AV2) Dataset and Evaluation Metrics
We further conduct long-range experiments on the recently released Argoverse 2 dataset [80] to demonstrate the superiority of our ContrastAlign in long-range detection. AV2 has a large scale dataset, and it contains 1000 sequences in total, 700 for training, 150 for validation, and 150 for testing. In addition to average precision (AP), AV2 adopts a composite score as an evaluation metric, which takes both AP and localization errors into account. The perception range in AV2 is 200 meters (cover area of 400m × 400m), which is much larger than Waymo. Such a large perception range leads to a huge memory footprint for dense detectors.
V-A3 Dataset and Evaluation Metrics.
We conducted experiments on the popular nuScenes[51], and Argoverse2 [80] datasets to validate the effectiveness of our approach. The detection ranges of the two datasets are 54 and 200 meters, respectively. For object detection on the nuScenes dataset, evaluation metrics include mAP and the nuScenes detection score (NDS). mAP is calculated by averaging over the distance thresholds of 0.5m, 1m, 2m, and 4m across all categories. NDS is a weighted average of mAP and five other true positive metrics that measure translation, scaling, orientation, velocity, and attribute errors. For object detection on the Argoverse2 dataset, mAP is adopted as the evaluation metric. We are testing on the AV2 dataset due to its extensive range of long-distance scenarios, where the issue of feature misalignment intensifies with increasing distance. In addition, to validate the robustness of feature alignment, we follow [52] in nuScenes dataset to simulate the misalignment between LiDAR and camera projection. It is worth noting that [52] only adds noise to the validation dataset, not to the training and testing datasets.
V-A4 Implementation Details.
We implement ContrastAlign within PyTorch [81], which is built upon the open-source OpenPCDet [82], where our baselines are BEVFusion [1] and GraphBEV [38]. In the LiDAR branch, feature encoding is performed by using TransFusion-L [28] to obtain LiDAR BEV features, where voxel dimensions are set to [0.075m, 0.075m, 0.2m] and point cloud ranges are specified as [-54m, -54m, -5m, 54m, 54m, 3m] along the X, Y, and Z axes, respectively. The camera branch employs a Swin Transformer [83] as the camera backbone, integrating heads of numbers 3, 6, 12, and 24, and utilizes FPN [84] for fusing multi-scale feature maps. The resolution of input images is adjusted and cropped to 256 $\times$ 704. In the LSS [3] configuration, frustum ranges are set with X coordinates [-54m, 54m, 0.3m], Y coordinates [-54m, 54m, 0.3m], Z coordinates [-10m, 10m, 20m], and depth ranges are set to [1m, 60m, 0.5m]. During training, we apply data augmentation for 10 epochs, which includes random flips, rotations (within the range $\frac{\pi}{4}$, $\frac{\pi}{4}$]), translations (with std=0.5), and scaling in the range of 0.9 to 1.1 for LiDAR data enhancement. We utilize CBGS [85] to resample the training data. Additionally, we use random rotation in $^{\circ}, 5.45.4^{\circ}$] and random resizing in [0.38, 0.55] to augment the images. The Adam optimizer [86] is used with a one-cycle learning rate policy, setting the maximum learning rate to 0.001 and weight decay to 0.01. The batch size is 24, and training is conducted on 8 NVIDIA GeForce RTX 3090 24G GPUs. During inference, we remove Test Time Augmentation (TTA) data augmentation, and the batch size is set to 1 on an A100 GPU. All latency measurements are taken on the same workstation with an A100 GPU.
V-B Main Results
Results on NuScenes Val set. As shown in Table I, we compare our ContrastAlign with SOTA methods on the nuScenes val set under clean and noisy misalignment settings. Compared to SOTA methods  [72, 28, 39], our method achieves the best performance, especially in challenging categories like construction vehicles, trailers, barriers, bikes, and traffic cones, which mainly consist of small and difficult-to-detect objects. On the clean nuScenes val set, our method achieves SOTA performance with 70.3% mAP and 72.5% NDS, outperforming the baseline BEVFusion  [1] by 1.8% mAP and 1.1% NDS, and achieving 71.5% mAP and 73.7% NDS in the baseline GraphBEV  [38], surpassing it by 1.4% mAP and 0.8% NDS.
Results on MisAlignment in NuScenes Val Set. As shown in Table II, our main focus is not on the clean benchmark but on addressing the noisy misalignment issue. In comparison to TransFusion  [28] and BEVFusion  [1] on the nuScenes val set under noisy misalignment setting, our ContrastAlign achieves SOTA performance. Compared to the baselines BEVFusion  [1] and GraphBEV  [38], our ContrastAlign improves performance by 7.3% and 14.1% mAP for spatial and temporal misalignment in BEVFusion, and by 1.4% and 11.1% mAP in GraphBEV. This demonstrates that our method significantly alleviates the misalignment issue.It is worth noting that although SOTAs like CMT  [73] perform well in spatial misalignment, they are not effective in solving temporal misalignment, which is because they only involve feature queries, while we consider the essence of instance alignment. It is worth noting that TransFusion  [28], as a feature-level method, addresses feature misalignment through attention. However, the global query scheme of TransFusion cannot be directly applied to BEVFusion  [1, 2] due to differences in their architectures or objectives. In summary, our ContrastAlign has shown significant improvement not only in the clean setting but also in the noise setting, proving th at the proposed method enables to deal with the misalignment issue.
Results on NuScenes Test Set. As shown in Table III, ContrastAlign achieves 72.9% mAP and 74.9% NDS on the nuScenes test benchmark, outperforming GraphBEV  [38] by 1.2% mAP and 1.3% NDS. And ContrastAlign outperforms BEVFusion [1] by by 2.0% mAP and 0.9% NDS. Compared to both LiDAR-based methods such as VoxelNeXt  [68], TransFusion-L  [28], and multi-modal methods such as TransFusion  [28], DeepInteraction  [30], and ObjectFusion  [39], our method achieves SOTA performance. Specifically, compared to ObjectFusion  [39], as a solution to BEV feature misalignment, our method outperforms ObjectFusion by 0.8% mAP and 0.5% NDS. What’s more, ContrastAlign has significant improvements in construction vehicles, barriers, motorcycles, bikes, pedestrians, and traffic cones, respectively. Although the nuScenes is an open-source clean dataset, it inevitably has minimal feature misalignment issues. Our notable improvements on the nuScenes test set indicate that our method effectively mitigates feature misalignment.
Results on Argoverse2 Val Set. To validate the effectiveness of our ContrastAlign on the long-range detection, we conducted experiments on the Argoverse2 dataset with a perception range of 200 meters as shown in Table IV. It is worth noting that the proposed ContrastAlign compared with the orther methods including CenterPoint [67], FSD [79], VoxelNeXt [68], BEVFusion [1] and GraphBEV [38], also outperforms the best. Specifically, our ContrastAlign achieves a 2.0% mAP improvement over GraphBEV, highlighting the method’s increased effectiveness with greater distances and more severe feature misalignment. This indicates that the farther the distance and the more severe the feature misalignment, the more effective ContrastAlign is.
V-C Robustness Analysis
To demonstrate the robustness of our method, we present the results in Table V under different times, ego distances, and object sizes. Since different times pose challenges to the detection performance of the model, we follow the procedure of BEVFusion  [1] and split the val set into Day/Night based on the keywords ‘day’ and ‘night’ in the scene descriptions. Compared to BEVFusion  [1], our method significantly improves Day and Night scenarios through contrastive learning for feature alignment. Specifically, for night scenarios, our method outperforms BEVFusion by 1.3%, where depth estimation presents a greater challenge for BEVFusion due to poor lighting conditions.
Far-small objects are more susceptible to feature misalignment. Following BEVFusion  [1], we categorize annotation and prediction ego distances into three groups, Near (0-20m), Middle (20-30m) and Far (>>30m), summarize the size distributions for each category, defining three equal-proportion size levels, Small, Moderate and Large. For the different ego distance, we have seen improvements in Near, Middle, and Far, especially in Far, which has increased by 3.9%. This indicates that the farther the distance, the more severe the feature misalignment and the more effective ContrastAlign is. In the different object sizes, we have seen significant improvements in Small, Moderate, and Large, especially in Small objects, which has increased by 3.1%. The improvement of the ‘Far’ and ‘Small’ metrics emphasizes the improvement of our method, which makes our method be visually meaningful since better feature alignment allows for better detection of smaller and farther objects.
V-D Ablation Studies
@@
Roles of different modules in ContrastAlign. To analyze the impact of different modules in ContrastAlign on feature misalignment, we conduct experiments on nuScenes validation set under the noise misalignment setting, as shown in Table VI. We have added C-Instance and L-Instance modules based on the baseline (BEVFusion), simply extracting features from them through additional detection heads such as CenterPoint head and VoxelNeXt head, and then adding them to existing BEV features. We extract RoI features of LiDAR and camera through C-Instance and L-Instance modules and add InstanceFusion to solve the misalignment problem through contrastive learning. It can be seen that after adding InstanceFusion module, there is a significant improvement, with mAP increasing from 61.2% to 68.1% (an increase of 6.9%), which is a very significant performance improvement. Meanwhile, the impact on latency is within a controllable range. Evidence demonstrates the effectiveness of incorporating contrastive learning into multi-modal 3D object detection to address feature misalignment issues in BEV features.
@@
Effect of the Hyperparameters $K$ and $\tau$ in InstanceFusion module. To assess the influence of various hyperparameters within the InstanceFusion module, an examination of the variables $\tau$ and $K$ was conducted on nuScenes val set under noisy misalignment setting. Where $\tau$ comes from Eq.(5), which is the threshold of IoU affecting the number of positive sample pairs, and $K$ comes from Eq.(7), affecting the number of negative samples. Due to the different positive and negative sample pairs at different frames, we cannot provide a fixed number. As shown in Table VII, when $\tau$ changes, $K$ is fixed at 8 to obtain better performance indicators. As shown in Table VIII, when $K$ changes, $\tau$ is fixed at 0.1. We choose several representative results for analysis. The larger the $\tau$, the smaller the number of positive samples. It can be observed that when $\tau$ is 0.1, the performance is the best, reaching 68.1% mAP and 70.9% NDS. The larger $K$, the larger the number of negative samples, and the best performance is achieved when $K$ is 8.
VI Conclusion